# LASER: Attention with Exponential Transformation

### v1

@@ 4.3 Vision Transformer (ViT) and Conformer - Speech-to-Text
In this section, we experiment with the Vision Transformer (ViT) (Dosovitskiy et al., 2021) variant - S/16 on the Imagenet-1k classification task (Deng et al., 2009) which is a part of AlgoPerf benchmarks (Dahl et al., 2023) for optimizer comparisons. These benchmarks are identically implemented in init2winit framework (Gilmer et al., 2023), build on JAX, which we use for our experiments. A hyperparameter sweep was conducted over 50 configurations on NAdamW (Dozat, 2016), focusing on the search space defined in Table 3. We selected the best-performing hyperparameter configuration based on validation performance for standard attention, run it for 5 different random seeds (for initialization) and report the validation curves corresponding to median in Figure 5, where we show that LASER attention provides a 1.15% absolute improvement in error rate (25.27% $\to$ 24.12%), i.e., a $\sim$4relative improvement over standard attention.
@@
We also evaluate the performance of LASER attention on the Librispeech Speech-to-Text dataset (Panayotov et al., 2015) using the Conformer model (Gulati et al., 2020). Similar to the ViT experiments, we use the AlgoPerf benchmark and perform a hyperparameter sweep across 50 configurations to optimize standard attention. We pick the optimal hyperparameters, run them for 5 different random seeds (for initialization) and report the validation curves corresponding to median in Figure 5 where we demonstrate a clear reduction in word error rate (WER) (0.0843 $\to$ 0.0824) when using LASER attention.
(a) Conformer Speech-to-Text - Validation
(b) ViT ImageNet Classification - Validation
Figure 5: Comparison of LASER attention vs Standard attention in two tasks: Conformer Speech-to-Text (left) and ViT ImageNet Classification (right). LASER attention provides a 1% absolute improvement in error rate (25.27% $\to$ 24.09%) i.e., a $\sim$4.67relative improvement. In Conformer, we notice an improvement of word error rate (WER) (0.0843 $\to$ 0.0824) - 2.25% relative improvement. Median curve among 5 random initializations corresponding to the best performing hyperparameter configuration is reported.
Comparisons using OPTLists. In AlgoPerf (Dahl et al., 2023), a list of 5 hyperparameters for NAdamW, tuned for a variety of benchmarks (as opposed to just Conformer and ViT) were provided - OPTList. We also evaluate our models on OPTList by running each hyperparameter with 5 different random seeds (initializations) and picking the median of the best performing hyperparameter. For Imagenet-ViT benchmark we obtain a reduction in validation error rate from 0.21732 to 0.21348. In Librispeech-Conformer benchmark we obtain a reduction in validation word error rate from 0.07728 to 0.07607.
5 Conclusions
@@
In this paper, we first identified a bottleneck in the gradient backpropagation of attention mechanism where the gradients are scaled by small Jacobian values while passing through the softmax operation. We fix this issue by transforming the inputs and outputs of attention mechanism, and show that this leads to larger Jacobians in the limiting case. We demonstrate the improvements in training performance over four types of Transformers spanning different modalities (text, speech and vision): (a) decoder-only (via Large Language model) upto 2.2 billion parameters, (b) encoder-only (BERT) with 2.2 billion parameters, (c)  vision Transformers on Imagenet, and (d) Conformer on Librispeech speech-to-text, where we show significant and consistent improvements in performance.
6 Limitations
Research on novel attention methods have wide applicability, however, due to the quadratic complexity in sequence length, scaling to large sequence lengths can be a major limitation of attention methods.

### v2

@@ 4.3 Vision Transformer (ViT) and Conformer - Speech-to-Text
In this section, we experiment with the Vision Transformer (ViT) (Dosovitskiy et al., 2021) variant - S/16 on the Imagenet-1k classification task (Deng et al., 2009) which is a part of AlgoPerf benchmarks (Dahl et al., 2023) for optimizer comparisons. These benchmarks are identically implemented in init2winit framework (Gilmer et al., 2023), build on JAX, which we use for our experiments. A hyperparameter sweep was conducted over 50 configurations on NAdamW (Dozat, 2016), focusing on the search space defined in Table 8. We selected the best-performing hyperparameter configuration based on validation performance for standard attention, ran it for 5 different random seeds (for initialization) and report the corresponding training runs in Table 4, where we show that LASER attention provides a 1.15% absolute improvement in error rate (25.32% $\to$ 24.17%), i.e., a $\sim$4relative improvement over standard attention.
Method Valid Error Test Error Train Error
Standard 0.2532 0.3749 0.1678
Standard + temp 0.2513 0.3742 0.1668
Standard + QK-normalization 0.2455 0.3644 0.1557
Standard + per-dim temp 0.2607 0.3821 0.1775
LASER 0.2417 0.3593 0.1527
LASER + temp 0.2427 0.3633 0.1538
LASER + QK-normalization 0.2372 0.3588 0.1427
LASER + per-dim temp 0.2475 0.3698 0.1573
Table 4: Error rates for ViT-S/16 on Imagenet. We bold the lowest errors within Standard and LASER attention variants. LASER maintains upto 1.15% improvement over standard attention variants.
Table 4 also highlights the consistent improvement of LASER over standard attention across different configurations. Temperature tuning does not introduce major changes in performance for either attention mechanism, though lower temperature values tend to degrade results due to sharper attention probability distributions. Adding QK-normalization substantially enhances performance, indicating its robustness across attention mechanisms. On the otherhand, adding per-dim temperature negatively impacts performance, likely due to hyperparameter tuning being optimized without its presence.
@@
We also evaluate the performance of LASER attention on the Librispeech Speech-to-Text dataset (Panayotov et al., 2015) using the Conformer model (Gulati et al., 2020). Similar to the ViT experiments, we use the AlgoPerf benchmark and perform a hyperparameter sweep across 50 configurations to optimize standard attention. We pick the optimal hyperparameters, run them for 5 different random seeds (for initialization) and report the validation curves corresponding to median in Table 5 where we demonstrate a reduction in word error rate (WER) (0.0828 $\to$ 0.0808) when using LASER attention.
Method Valid WER Test WER Train WER
Standard 0.083192 0.050876 0.032594
Standard + temp 0.08362 0.050398 0.032275
Standard + QK-Normalization 0.082865 0.049747 0.029574
LASER 0.082892 0.049786 0.030084
LASER + temp 0.082137 0.049384 0.032822
LASER + QK-Normalization 0.080837 0.049231 0.031283
Table 5: Word Error Rate (WER) results for Conformer model with different attention mechanisms and modifications. We bold the lowest WER values within Standard and LASER attention variants. LASER shows upto 0.2% improvements on word error rates.
In Table 5, we found adding QK-Normalization enhances performance for both LASER and standard attention. Temperature tuning does not significantly impact standard attention but provides slight benefits to LASER . Since per-dim temperature is built in Conformer model architecture (Dahl et al., 2023) we don’t conduct an ablation study on its effect.
Ablations in Large Language Modeling
In this section, we apply LASER and observe how it performs under the following changes: a) increase the model parameters to test the scalability of LASER b) using a Diff Transformer (Ye et al., 2024).
Scalability of LASER
To test the scalability of LASER, we trained a model with 7.7B parameters for 44B tokens. Compared to the 2.2B parameter model (which had model dimension 2048, hidden dimension 8192, 32 layers, and 8 attention heads each of size 512), the 7.7B parameter model was scaled along all dimensions: model dimension 3440, hidden dimension 11584, number of heads 16, and dimension per head 720.
The downstream task evaluation for the 7.7B parameter models is presented in Table 6.
Table 6: Downstream task accuracies for 7.7B parameter models trained on 44B tokens. LASER shows an average improvement of 1.44% over standard attention. More pronounced differences are observed in tasks like BoolQ (+6%), CB (+1.78%), OpenbookQA (+1.8%), and RTE (+3.61%).
Dataset LASER Standard
ArcE 52.48 52.69
BoolQ 62.45 56.48
CB 44.64 42.86
HellaSwag 57.16 56.02
MultiRC 56.00 55.59
OpenbookQA 45.40 43.60
RaceM 44.64 44.15
RTE 53.43 49.82
StoryCloze 71.78 71.51
WiC 47.34 47.34
Winogrande 58.41 57.77
Average 53.97 52.53
4.4 Ablation with Diff Transformer
The Differential Transformer (DiffTransformer) (Ye et al., 2024) introduces a differential attention mechanism. This mechanism calculates attention scores as the difference between two separate softmax attention maps to cancel noise and promote sparse attention patterns (Ye et al., 2024). The Diff Transformer attention mechanism is formulated as:
$\text{DiffAttn}(X)$ $=\text{softmax}\left(\frac{Q_{1}K_{1}^{T}}{\sqrt{d_{k}}}\right)V$
$-\lambda\text{softmax}\left(\frac{Q_{2}K_{2}^{T}}{\sqrt{d_{k}}}\right)V$
where $Q_{1},Q_{2},K_{1},K_{2}\in\mathbb{R}^{N\times d_{k}}$ are projected query and key vectors, $V\in\mathbb{R}^{N\times d_{v}}$ is the projected value vector (in the original DiffTransformer paper, $d_{v}=2d_{k}$ for the combined $V$), $\lambda$ is a learnable scalar, and $d_{k}$ is the dimension of the key/query vectors per head.
We propose a modification, LASER+DiffTransformer, by equipping the softmax dot-product attention terms with LASER modifications. The LASER attention mechanism involves conducting attention on exponentially transformed inputs, $\exp(V)$, and takes the logarithm of the result. Applying this to DiffTransformer, we get:
$\text{DiffAttn}_{\text{LASER}}(X)$ $=\log\left(\text{softmax}\left(\frac{Q_{1}K_{1}^{T}}{\sqrt{d_{k}}}\right)\exp(V)\right)$
$-\lambda\log\left(\text{softmax}\left(\frac{Q_{2}K_{2}^{T}}{\sqrt{d_{k}}}\right)\exp(V)\right)$
This formulation aims to address potential backpropagation challenges due to the softmax operations in both attention maps of the DiffTransformer. We trained a 2.2B parameter DiffTransformer with Model dim: 2048, Hidden dim: 8192, Number of attention heads: 8, Head size: 512, and Number of layers: 32. The models were trained on 24 billion tokens. The results are shown in Table 7.
Table 7: Downstream task performance for 2.2B parameter DiffTransformer and DiffTransformer+LASER models, trained on 24 billion tokens. On average, Diff+LASER shows an improvement of approximately 1% over DiffTransformer.
Dataset Diff+LASER DiffTransformer
ArcE 49.28 49.20
CB 42.85 41.07
HellaSwag 51.93 51.58
MultiRC 55.13 52.82
OpenbookQA 44.20 43.00
RaceM 42.40 40.73
RTE 52.34 50.90
StoryCloze 71.03 71.29
WiC 50.00 49.53
Winogrande 56.04 55.80
Average 51.52 50.59
5 Conclusion
@@
We identified a bottleneck in the gradient backpropagation of attention mechanism where the gradients are scaled by small Jacobian values while passing through the softmax operation. We alleviate this issue by transforming the inputs and outputs of attention mechanism, and show that this leads to larger Jacobians in the limiting case. We demonstrate the improvements in training performance over four types of Transformers spanning different modalities (text, speech and vision): (a) decoder-only (via Large Language model) upto 2.2 billion parameters, (b)  vision Transformers on Imagenet, (c) Conformer on Librispeech speech-to-text and (d) encoder-only (BERT) with 2.2 billion parameters,, where we show significant and consistent improvements in performance.
Acknowledgements
We thank Devvrit Khatri, Nilesh Gupta and Aditya Kusupati for valuable suggestions.
Impact Statement
The paper presents a technique which improves pretraining of transformers, a widely used architecture in machine learning. There are many positive potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.2 Proofs
A.3 Training loss curves
(a) Train loss - 2.2 billion parameter model.
(b) Test loss - 2.2 billion parameter model.
(c) Train loss - 1.1 billion parameter model.
(d) Test loss - 1.1 billion parameter model.
Figure 5: Performance comparison for 2.2 billion and 1.1 billion parameter models. The 2.2 billion model has 32 layers, 8 attention heads with head size 512, MLP hidden dimension 8192, and model dimension 2048. The 1.1 billion model has 32 layers, 8 attention heads (head size 256), MLP hidden dimension 4096, and model dimension 1024. LASER demonstrates better train and test loss compared to Standard Attention even in large-scale settings.
A.4 Fluctuations in downstream evaluations
(a) Part 1 Dataset LASER Standard WSC 80.98$\pm$0.68 79.30$\pm$0.44 Winogrande 62.19$\pm$0.21 62.15$\pm$0.19 Winograd 82.27$\pm$0.37 80.29$\pm$0.27 WiC 51.44$\pm$0.66 51.29$\pm$0.55 StoryCloze 77.95$\pm$0.06 76.46$\pm$0.07 RTE 53.14$\pm$0.14 53.14$\pm$0.14 ReCoRD 85.24$\pm$0.04 85.06$\pm$0.05 RaceM 50.56$\pm$0.18 49.67$\pm$0.13
(b) Part 2 Dataset LASER Standard RaceH 37.99$\pm$0.13 37.56$\pm$0.11 PIQA 77.20$\pm$0.12 76.74$\pm$0.12 OpenBookQA 49.08$\pm$0.20 47.56$\pm$0.15 MultiRC 57.16$\pm$0.18 53.99$\pm$0.16 HellaSwag 66.61$\pm$0.07 65.46$\pm$0.04 COPA 82.00$\pm$0.00 80.00$\pm$0.00 CB 40.00$\pm$0.87 43.93$\pm$0.87 BoolQ 63.39$\pm$0.18 60.45$\pm$0.38
Table 10: Here we add standard deviations of each number, computed on downstream evaluations of 5 checkpoints and note an average for LASER to be 63.58$\pm$0.26 and an average for standard attention to be 62.69$\pm$0.23, noting a difference of 0.89.
A.5 Training Analysis
Training instability. There can be spikes in training loss curves initially during large language model training. We notice that despite these spikes training stabilizes and converges smoothly. However, training instability/spikes can be attributed to poor model architecture and optimizer choices. We now ablate the choice of attention mechanism and understand its affect on training stability. Figure 6 compares training stability of different models.
(a) 234M Parameters
(b) 300M Parameters
(c) 1.1B Parameters
(d) 2.2B Parameters
Figure 6: Train loss vs. steps for LASER and standard attention across different parameter scales. Training stability for each attention mechanism can be observed through the number of training spikes. Generally, models with LASER attention exhibit fewer training spikes compared to models with standard attention, indicating greater stability in training for LASER attention across all parameter scales. The figures focus on the initial part of the training as the remainder did not exhibit training instability.
Additionally, we ran pretraining of a 2.2B parameter language model presented in Section 4.1 with BF16 precision weights and activations and compare LASER and Standard attention in Table 11.
Precision LASER Loss Standard Attention Loss
FP32 2.326 2.344
BF16 2.333 2.350
Table 11: Comparison of LASER and Standard Attention Losses for FP32 and BF16 precision on a 2.2B model.
We cached 4800 query, key, value matrices of size (1024, 8, 256) (headsize-256, heads-8, sequence length-1024) during training, and computed the following numerical reconstruction relative errors of attention output in bfloat16: vanilla LASER (0.0018, 0.0002), LASER+log-weighted-sum-exp-trick (0.0017,0.0001), Standard Attention (0.0016, 0.0001). This experiment is conducted in a machine with 4 TPUv5 chips. We conducted the same experiment on 16 A100s and found the following errors: vanilla LASER (0.002, 0.0003), LASER+log-weighted-sum-exp-trick (0.0019,0.0002), Standard Attention (0.0018, 0.0002). While on an average, we found log-weighted-sum-exp-trick to help on both TPUv5 and A100s, we note that this trick prevents overflows, which is crucial for stable training. Similar trick famously known as log-sum-exp is used to prevent overflows due to exp(.) function in the softmax function and is adopted by both Pytorch and Jax in their softmax implementations.
Understanding training instability. The enhanced training stability of the LASER attention mechanism can be formally attributed to its direct solution for the gradient saturation problem inherent in standard attention models. The Jacobian of the softmax function, which scales the gradients for the query and key projection matrices ($W_{Q}$ and $W_{K}$), diminishes towards zero when attention probabilities become sharp. This vanishing gradient signal create inconsistent updates to parameters, leading to suboptimal learning and potential loss spikes. LASER circumvents this by reformulating the attention output as $\log(\text{softmax}(QK^{T})\exp(V))$ , a structure which admits with lower saturation properties.
This improved parameter learning grants LASER its stability, a behavior analogous to residual connections in deep networks. The LASER output for a given token $i$, $o_{i}$, can be expressed as a log-sum-exp function, which is a tight, differentiable approximation of the $\max$ function. Here we bound $o_{1}$:
$\max$ $(v_{1}+\log(a_{11}),v_{2}+\log(a_{12}))$
$\leq o_{1}\leq\max(v_{1}+\log(a_{11}),v_{2}+\log(a_{12}))+\log(2).$
In cases of strong self-attention, where the attention probability $a_{ii}$ is high, the $\max$ function property ensures that the output $o_{i}$ closely approximates its corresponding input value $v_{i}$. By creating this effective ”pass-through” for the input signal during moments of high certainty, LASER functions as an inherent residual connection. We conjecture that this property is the formal mechanism that underpins the empirically observed reduction in training loss spikes compared to standard attention.